# On Three-Layer Data Markets

### v2

@@ 6 Regulations
  • Otherwise, if there is at least one high-cost platform, then there exists $\tilde{\beta}$, $\underline{\beta}$, and $\bar{\sigma}$ such that
@@ 7 Conclusion
In this paper, we study the intricate interplay between user privacy concerns and the dual incentives of online platforms: to safeguard user privacy and to monetize user data through sales to third-party buyers. To this end, we have constructed a multi-stage game model that captures the interactions between users, platforms, and data buyers. Our model delineates various scenarios where the interplay of platform service costs, the valuation of data, and privacy concerns dictate equilibrium outcomes. Additionally, we examine potential regulations designed to bolster user privacy and enhance market outcomes for users.
There are several exciting avenues for future research. In our framework, we have modeled a user as a representative of a market segment. Extending our analysis to a setting with multiple heterogeneous users is very interesting. This is because data externalities among different users may cause indirect information disclosure, which in turn could alter both user strategies and potential regulations aimed at enhancing user experience. Similarly, our focus has been on a representative data buyer who aims to acquire user data and learn their preferences. Extending our findings beyond this is very valuable. This is because user data is inherently multi-dimensional and can be aggregated from diverse sources, potentially modifying user strategies and calling for tailored regulations. Lastly, we have investigated regulations that restrict the transfer of user data to buyers. The challenge of developing alternative regulatory frameworks that address the intricacies of the data market—frameworks that are both adaptable and resilient enough to keep pace with swift technological advancements—remains an open and important area for exploration.
8 Acknowledgment
Alireza Fallah acknowledges support from the European Research Council Synergy Program, the National Science Foundation under grant number DMS-1928930, and the Alfred P. Sloan Foundation under grant G-2021-16778. The latter two grants correspond to his residency at the Simons Laufer Mathematical Sciences Institute (formerly known as MSRI) in Berkeley, California, during the Fall 2023 semester. Michael Jordan acknowledges support from the Mathematical Data Science program of the Office of Naval Research under grant number N00014-21-1-2840 and the European Research Council Synergy Program.
Appendix A Proofs:
This appendix included the omitted proofs from the text.
@@ Appendix A Proofs:
Without loss of generality, we prove the claim for $S=[K]$. Let us recall a property of normal distributions: If $\mathbf{x}$ is a $n$-dimensional normal random variable is partitioned as
${\mathbf{x}}={\begin{bmatrix}{\mathbf{x}}_{1}\\ {\mathbf{x}}_{2}\end{bmatrix}}{\text{ with sizes }}{\begin{bmatrix}k\times 1\\ (n-k)\times 1\end{bmatrix}}$
with with sizes
$\begin{bmatrix}k\times k&k\times(n-k)\\ (n-k)\times k&(n-k)\times(n-k)\end{bmatrix}$
and accordingly $\bm{\mu}$ is partitioned as
$\bm{\mu}=\begin{bmatrix}\bm{\mu}_{1}\\ \bm{\mu}_{2}\end{bmatrix}$
with sizes
$\begin{bmatrix}k\times 1\\ (n-k)\times 1\end{bmatrix}$
and $\bm{\Sigma}$ is partitioned as
$\bm{\Sigma}=\begin{bmatrix}\bm{\Sigma}_{11}&\bm{\Sigma}_{12}\\ \bm{\Sigma}_{21}&\bm{\Sigma}_{22}\end{bmatrix}$
with sizes
$\begin{bmatrix}k\times k&k\times(n-k)\\ (n-k)\times k&(n-k)\times(n-k)\end{bmatrix}$
then the distribution of $\mathbf{x}_{1}$ conditional on $\mathbf{x}_{2}=\mathbf{a}$ is normal with covariance matrix
@@
The proof of this lemma follows directly from the above fact. $\blacksquare$
@@
Both parts follow from the conditional expectation being the MSE estimator. To see the first part, notice that we can always throw away data points. To see the second part, notice that we can always add extra noise to the data points. $\blacksquare$
@@
We first argue that, for any $i\in[K]$, the equilibrium price of platform $i$ must be such that the buyer finds it optimal to purchase from this platform in the corresponding buyer equilibrium. To see this notice that if the buyer does not purchase from platform $i$, the platform’s utility from interacting with the buyer becomes 00. Now, if the platform deviates and selects the price
@@
then the buyer always finds it optimal to purchase from this platform, and therefore, the platform’s utility weakly increases from 00. This is because for any $\bm{b}_{-i}$, we have
@@
$=\beta\left[\left(\mathcal{I}(\bm{\sigma},\bm{e},\bm{a},(b_{i}=1,\bm{b}_{-i}))-\mathcal{I}(\bm{\sigma},\bm{e},\bm{a},(b_{i}=0,\bm{b}_{-i}))\right)-\left(\mathcal{I}(\bm{\sigma},\bm{e},\bm{a},\mathbf{1})-\mathcal{I}(\bm{\sigma},\bm{e},\bm{a},(b_{i}=0,\mathbf{1}_{-i}))\right)\right]$
$\overset{(a)}{\geq}0$
where (a) follows from Lemma 3. $\blacksquare$
@@
Suppose the contrary that there exists a platform equilibrium strategy $(\bm{\sigma},\bm{e})$ for which $e_{i}=1$, $a_{i}=0$ and $\bm{a}\in\mathcal{A}^{\mathrm{E}}(\bm{\sigma},\bm{e})$. We argue that platform $i$ has a profitable deviation. In particular, notice that the current utility of platform $i$ is $-c_{i}$, as the buyer will not pay any non-zero price for this platform’s data. Now, if platform $i$ chooses $\tilde{e}_{i}=1$, $\tilde{p}_{i}=0$ and $\tilde{\sigma}_{i}=\infty$, then in any $\bm{a}\in\mathcal{A}^{\mathrm{E}}((\tilde{\sigma}_{i},\bm{\sigma}_{-i}),\bm{e})$ we must have that $a_{i}=1$ and therefore the platform $i$’s utility increases to $\mathcal{I}_{i}-c_{i}$. $\blacksquare$
@@
$\mathcal{U}_{1}((\sigma_{1},\sigma_{2})),\mathbf{1},\mathbf{p}^{*},\mathbf{1})=\frac{1}{2}+\beta\left(\mathcal{I}(\sigma_{1},\sigma_{2})-\mathcal{I}(\sigma_{2})\right)\geq\frac{1}{2}+\beta\mathcal{I}(\tilde{\sigma}_{1})=\mathcal{U}_{1}((\tilde{\sigma}_{1},\sigma_{2})),(1,0)),\mathbf{p}^{*},\mathbf{1}).$ (15)
@@
and thus, $\mathcal{I}(\sigma_{1},\sigma_{2})=1/\alpha$. Using this, we can rewrite the condition (15) as
@@
$1-\alpha\mathcal{I}(\tilde{\sigma_{1}},\sigma_{2}^{A})=\frac{1}{2}-\alpha\mathcal{I}(\tilde{\sigma_{1}}).$ (21)
@@
We can verify that the right-hand side minus the left-hand side is equal to zero at $\bar{\alpha}\approx 1.88$ and positive afterward. Notice that $\eta$ and therefore the condition (24) is symmetric with respect to both platforms, and consequently, considering the other deviation to $(\sigma_{1}^{A},\tilde{\sigma}_{2})$ leads to the exact same condition on $\alpha$. This completes the proof of our claim that either both conditions hold if $\alpha\geq\bar{\alpha}$ and neither hold if $\alpha<\bar{\alpha}$. $\blacksquare$
@@
Proof of Lemma 7: Using Proposition 2, we know that the equilibrium noise variance belongs to $\Sigma_{\mathbf{1},\bm{e}}$. We next argue that it cannot belong to the interior of $\Sigma_{\mathbf{1},\bm{e}}$. To see this, suppose the contrary that $\bm{\sigma}\in\mathrm{int}(\Sigma_{\mathbf{1}},\bm{e})$ is an equilibrium. There exists $i\in[K]$ such that by locally decreasing $\sigma_{i}$, we remain in the set $\Sigma_{\mathbf{1},\bm{e}}$. We claim that platform $i$ has an incentive to decrease its noise variance. First, note that by decreasing its noise variance, we remain in the set $\Sigma_{\mathbf{1},\bm{e}}$ and that the user still finds it optimal to share with platform $i$. However, the payment she receives from the buyer, which is given by
@@
$\frac{1}{2}-c_{i}+\beta\left(\mathcal{I}(\bm{\sigma},\bm{e},(a_{i}=1,\bm{a}_{-i}),\mathbf{1})-\mathcal{I}(\bm{\sigma},\bm{e},(a_{i}=0,\bm{a}_{-i}),\mathbf{1})\right)\geq 0.$
@@
Notice that, for any given $\beta_{-i}$, the denominator is positive because the amount of leaked information strictly decreases when one less user shares information. Finally, $\bar{\beta}$ is finite becasue the outer maximum is taken over finitely many elements.
@@
We next prove that assuming a set of platforms that have entered the market, for large enough $\alpha$, there exists an equilibrium. Without loss of generality, let us assume all platforms have entered the market. $\bm{\sigma}\in\Sigma_{\mathbf{1},\mathbf{1}}\cap\Sigma_{\mathbf{0},\mathbf{1}}$ is an equilibrium noise level if no platform has a deviation. Let us consider platform 1. For this platform, the possible deviations are to $\Sigma_{\mathbf{1},\mathbf{1}}\cap\Sigma_{\bm{a},\mathbf{1}}$ where $a_{1}=0$ and $\bm{a}_{-1}$ can be any vector in $\{0,1\}^{K-1}$. We next argue that $\bm{\sigma}\in\Sigma_{\mathbf{1},\mathbf{1}}\cap\Sigma_{\mathbf{0},\mathbf{1}}$ for which
@@
for any $S\subseteq[K]$ denote the leaked information about $\theta$ when the data of the user is shared with the buyer through platforms in the set $S$ and platform $i\in S$ is using noise variance $\sigma_{i}^{2}$.
@@
=$\overset{(a)}{=}$ $\bm{\gamma}^{T}\left(\bm{\gamma}\bm{\gamma}^{T}+(\frac{1}{t}-1)\mathrm{diag}(\gamma_{1}^{2},\dots,\gamma_{K}^{2})\right)^{-1}\bm{\gamma}$
=$\overset{(a)}{=}$ $\bm{\gamma}^{T}\left(\frac{1}{\frac{1}{t}-1}\mathrm{diag}(1/\gamma_{1}^{2},\dots,1/\gamma_{K}^{2})-\frac{\frac{1}{\frac{1}{t}-1}\mathrm{diag}(1/\gamma_{1}^{2},\dots,1/\gamma_{K}^{2})\bm{\gamma}^{T}\bm{\gamma}\frac{1}{\frac{1}{t}-1}\mathrm{diag}(1/\gamma_{1}^{2},\dots,1/\gamma_{K}^{2})}{1+\bm{\gamma}^{T}\frac{1}{\frac{1}{t}-1}\mathrm{diag}(\gamma_{1}^{2},\dots,\gamma_{K}^{2})\bm{\gamma}}\right)\bm{\gamma}$
@@
Let us consider the deviation of platform 1 so that the corresponding user equilibrium is $(a_{1}=1,\bm{a}_{-i})$ with $\bm{a}_{-i}$ having $i-1$ ones and $K-i$ zeros. Notice that if such deviation is not possible by increasing $\sigma_{1}$, then this is not a profitable deviation for platform 1, and we now verify that even if such deviation is possible, it is not profitable for platform 1. Given that $\bm{\sigma}\in\Sigma_{\mathbf{1},\mathbf{1}}\cap\Sigma_{\mathbf{0},\mathbf{1}}$, we have that the user is indifferent between sharing with all platforms and sharing with none of them. Therefore, we have
@@
Using Lemma 9, the above equality gives us
@@
Given the above deviation to $(\tilde{\sigma}_{1},\dots,\sigma_{K})$, the user indifference condition gives us
$\frac{K}{2}-\alpha\mathcal{I}(\tilde{\sigma}_{1},\dots,\sigma_{K})=\frac{i}{2}-\alpha\mathcal{I}(\tilde{\sigma}_{1},\dots,\sigma_{i}).$
Substituting the closed-form expressions of Lemma 9 in the above equality gives us
$t^{\prime}=\frac{2-4\alpha+K-i-\sqrt{16\alpha^{2}+(K-i)^{2}-8\alpha K}}{2+4\alpha-2K},$
@@
Now, for this deviation to be not profitable, we must have
@@
Again, using the closed-form expressions of Lemma 9 and writing $t^{\prime}$ and $t$ in terms of $\alpha$, the above inequality becomes equivalent to
@@
$\frac{3K-i}{4\alpha}+\frac{1-K}{2\alpha-1}+\frac{i-1}{2\alpha-1+i-K}-1+\frac{\sqrt{16\alpha^{2}+(K-i)^{2}-8\alpha K}}{4\alpha}$
@@
Notice that for any other set $S\subseteq[K]$ of the platforms that have entered the market, the above bound on $\alpha$ guarantees the existence of platform equilibrium strategy as $|S|\leq K$. This completes the proof. $\blacksquare$
@@
Clearly, there is no uniform minimum privacy mandate that can make the user better off because banning data sharing maximizes the user gain from services and minimizes the privacy loss.
@@
[…]
For any $\alpha$, as $\bar{\sigma}$ increases, the leaked information goes to zero, and therefore, so long as there exists at least one high-cost platform, a $\bar{\sigma}$ for which the above inequality holds exist. We next argue that all platforms enter the market for large enough $\beta$. To see this, we need to have
@@
for all $i\in[K]$. Therefore, by letting $\beta\geq\tilde{\beta}$ where
@@
Similar to the proof of Proposition 7 and Theorem 1, we choose $\bar{\alpha}$ and $\hat{\beta}$ large enough so that an equilibrium exists and all platforms enter the market. Now suppose $\bar{\sigma}$ is the optimal uniform minimum privacy mandate so that all platforms enter the market, the user utility is

### v3

@@ 6 Regulation
  • Otherwise, if there is at least one high-cost platform, then there exist $\tilde{\beta}$, $\underline{\beta}$, and $\bar{\sigma}$ such that
@@ 7 Conclusion
We have studied the intricate interplay between user privacy concerns and the dual incentives of online platforms: to safeguard user privacy and to monetize user data through sales to third-party buyers. To this end, we have constructed a multi-stage game model that captures the interactions between users, platforms, and data buyers. Our model delineates various scenarios where the interplay of platform service costs, the valuation of data, and privacy concerns dictate equilibrium outcomes. Additionally, we examine potential regulations designed to bolster user privacy and enhance market outcomes for users.
There are several exciting avenues for future research. In our framework, we have modeled a user as a representative of a market segment. Extending our analysis to a setting with multiple heterogeneous users would be very worthwhile. This is because data externalities among different users may cause indirect information disclosure, which in turn could alter both user strategies and potential regulations aimed at enhancing user experience. Similarly, our focus has been on a representative data buyer who aims to acquire user data and learn their preferences. It would be useful to consider an elaborated model with multiple buyers, given that user data is inherently multi-dimensional and can be aggregated from diverse sources, potentially modifying user strategies and calling for tailored regulation. Lastly, we have investigated regulation that restricts the transfer of user data to buyers. The challenge of developing alternative regulatory frameworks that address the intricacies of the data market—frameworks that are both adaptable and resilient enough to keep pace with swift technological advancements—will remain of significant importance for quite some timem to come.
Appendix A Missing Proofs
This appendix includes the omitted proofs from the text.
@@ Appendix A Missing Proofs
Without loss of generality, we prove the claim for $S=[K]$. We first recall a property of the multivariate normal distribution: If $\mathbf{x}$ is an $n$-dimensional normal random variable partitioned as
$\mathbf{x}=\begin{bmatrix}\mathbf{x}_{1}\\[2.0pt] \mathbf{x}_{2}\end{bmatrix}$,
with corresponding mean and covariance matrices
$\bm{\mu}=\begin{bmatrix}\bm{\mu}_{1}\\[2.0pt] \bm{\mu}_{2}\end{bmatrix},\quad\bm{\Sigma}=\begin{bmatrix}\bm{\Sigma}_{11}&\bm{\Sigma}_{12}\\[2.0pt] \bm{\Sigma}_{21}&\bm{\Sigma}_{22}\end{bmatrix}$,
then the conditional distribution of $\mathbf{x}_{1}\mid\mathbf{x}_{2}=\mathbf{a}$ is normal with covariance matrix
@@
Applying this result to our setting, let $x_{1}=\mathbf{y}^{\top}\bm{\theta}$ and $\mathbf{x}_{2}=(\tilde{\mathbf{S}}_{i}:i\in S)$, where each $\tilde{\mathbf{S}}_{i}=\mathbf{x}_{i}^{\top}\bm{\theta}+\mathcal{N}(0,1)+\mathcal{N}(0,\sigma_{i}^{2})$. Then:
  • $\Sigma_{11}$ is the variance of $\mathbf{y}^{\top}\bm{\theta}$, which equals $1$.
  • $\Sigma_{22}$ is a $|S|\times|S|$ matrix with diagonal entries $\mathrm{Var}(\tilde{\mathbf{S}}_{i})=2+\sigma_{i}^{2}$, and off-diagonal entries $\mathrm{Cov}(\tilde{\mathbf{S}}_{i},\tilde{\mathbf{S}}_{j})=\mathbf{x}_{i}^{\top}\mathbf{x}_{j}=\gamma_{i}\gamma_{j}$ by Assumption 1.
  • $\Sigma_{12}$ is a $1\times|S|$ vector whose $i$-th entry is $\mathrm{Cov}(\mathbf{y}^{\top}\bm{\theta},\tilde{\mathbf{S}}_{i})=\mathbf{y}^{\top}\mathbf{x}_{i}=\gamma_{i}$.
Therefore, the posterior variance of $Z=\mathbf{y}^{\top}\bm{\theta}$ conditional on $(\tilde{\mathbf{S}}_{i}:i\in S)$ is
$\mathrm{Var}(Z\mid(\tilde{\mathbf{S}}_{i}:i\in S))=1-\mathbf{m}^{\top}M^{-1}\mathbf{m},$
where $\mathbf{m}=(\gamma_{i})_{i\in S}$ and $M$ is the $|S|\times|S|$ matrix with $M]_{ii}=2+\sigma_{i}^{2}$ and $M]_{ij}=\gamma_{i}\gamma_{j}$ for $i\neq j$.
Thus, by definition, the revealed information becomes
$\mathcal{I}(\bm{\sigma}_{S})$ $=\mathbb{E}\left[\left(Z-\mathbb{E}_{Z\sim\pi}\left[Z\right]\right)^{2}\right]-\mathbb{E}\left[\left(Z-\mathbb{E}_{Z\sim\pi_{\mathcal{S}}}\left[Z\right]\right)^{2}\right]$
$=\mathrm{Var}(Z)-\mathrm{Var}(Z\mid(\tilde{\mathbf{S}}_{i}:i\in S))$
$=1-\left(1-\mathbf{m}^{\top}M^{-1}\mathbf{m}\right)=\mathbf{m}^{\top}M^{-1}\mathbf{m},$
which completes the proof. $\blacksquare$
@@
We make use of Lemma 1 in this proof. Throughout, $M$ is a covariance matrix defined in that lemma.
Monotonicity in the set $S$. Let $T=S\cup\{k\}$ where $k\notin S$. We use the subscripts $S$ and $T$ to indicate the dependence on the set and write the block decomposition.
$M_{T}=\begin{bmatrix}M_{S}&&&\bm{u}\\[2.0pt] \bm{u}^{\top}&&&a\end{bmatrix},\qquad\bm{m}_{T}=\begin{bmatrix}\bm{m}_{S}\\ m_{k}\end{bmatrix}$,
where $\bm{u}=(\gamma_{i}\gamma_{k})_{i\in S}$ and $a=2+\sigma_{k}^{2}$. Since $M_{S}\succ 0$, the Schur complement
$\Delta\;\equiv\;a-\bm{u}^{\top}M_{S}^{-1}\bm{u}\;>\;0.$
The block matrix inversion formula yields
$M_{T}^{-1}=\begin{bmatrix}M_{S}^{-1}+M_{S}^{-1}\bm{u}\,\Delta^{-1}\bm{u}^{\top}M_{S}^{-1}&-M_{S}^{-1}\bm{u}\,\Delta^{-1}\\[2.0pt] -\Delta^{-1}\bm{u}^{\top}M_{S}^{-1}&\Delta^{-1}\end{bmatrix}$.
Therefore, we can write
$\mathcal{I}(\bm{\sigma}_{T})=\bm{m}_{T}^{\top}M_{T}^{-1}\bm{m}_{T}=\bm{m}_{S}^{\top}M_{S}^{-1}\bm{m}_{S}+\big(m_{k}-\bm{u}^{\top}M_{S}^{-1}\bm{m}_{S}\big)^{2}\,\Delta^{-1}.$
Since $\Delta^{-1}>0$, the second term is nonnegative, implying
$\mathcal{I}(\bm{\sigma}_{T})\;-\;\mathcal{I}(\bm{\sigma}_{S})=\frac{\big(m_{k}-\bm{u}^{\top}M_{S}^{-1}\bm{m}_{S}\big)^{2}}{\,a-\bm{u}^{\top}M_{S}^{-1}\bm{u}\,}\;\geq 0.$
Therefore, $\mathcal{I}(\bm{\sigma}_{S})$ is increasing in $S$.
Monotonicity in $\bm{\sigma}_{S}$. Let us denote $\mathcal{I}(\bm{\sigma}_{S})=\mathbf{m}^{T}M^{-1}\mathbf{m}$ with $M=M(\bm{\sigma}_{S})$ depending on $\sigma_{i}$ only through the diagonal entries. Therefore, we have $\partial M/\partial\sigma_{i}=2\sigma_{i}\bm{e}_{i}\bm{e}_{i}^{\top}$, where $\bm{e}_{i}$ is the $i$th standard basis vector (for $i\in S$). Using the matrix derivative identity
$\frac{\partial M^{-1}}{\partial\sigma_{i}}=-M^{-1}\!\left(\frac{\partial M}{\partial\sigma_{i}}\right)\!M^{-1},$
we obtain
$\frac{\partial\mathcal{I}(\bm{\sigma}_{S})}{\partial\sigma_{i}}$ $=\bm{m}^{\top}\!\left(\frac{\partial M^{-1}}{\partial\sigma_{i}}\right)\!\bm{m}=-\bm{m}^{\top}M^{-1}\!\left(2\sigma_{i}\,\bm{e}_{i}\bm{e}_{i}^{\top}\right)\!M^{-1}\bm{m}$
$=-2\sigma_{i}\,\big(\bm{e}_{i}^{\top}M^{-1}\bm{m}\big)^{2}\;\leq\;0.$
Therefore, $\mathcal{I}(\bm{\sigma}_{S})$ is decreasing in each $\sigma_{i}$. $\blacksquare$
@@
We first argue that, for any $i\in[K]$ with $a_{i}=e_{i}=1$, the equilibrium price of platform $i$ must be such that the buyer finds it optimal to purchase from this platform in the corresponding buyer equilibrium. To see this notice that if the buyer does not purchase from platform $i$, the platform’s utility from interacting with the buyer becomes 0. Now, if the platform deviates and selects the price
@@
then the buyer always finds it optimal to purchase from this platform, and therefore, the platform’s utility weakly increases from 0. This is because for any $\bm{b}_{-i}$, we have
@@
$=\beta\left[\left(\mathcal{I}(\bm{\sigma},\bm{e},\bm{a},(b_{i}=1,\bm{b}_{-i}))-\mathcal{I}(\bm{\sigma},\bm{e},\bm{a},(b_{i}=0,\bm{b}_{-i}))\right)\right$.
$\quad\left.-\left(\mathcal{I}(\bm{\sigma},\bm{e},\bm{a},\mathbf{1})-\mathcal{I}(\bm{\sigma},\bm{e},\bm{a},(b_{i}=0,\mathbf{1}_{-i}))\right)\right]\overset{(a)}{\geq}0$
where (a) follows from Lemma 3. This establishes that, at equilibrium, the buyer will purchase data from all platforms that have entered the market and possess user data.
Now, suppose $\bm{p}$ is a price equilibrium. Fix $i$. On the one hand, the buyer has no incentive to deviate and refrain from buying data from the platform $i$ if and only if
$p_{i}\leq\beta\left(\mathcal{I}(\bm{\sigma},\bm{e},\bm{a},\mathbf{1})-\mathcal{I}(\bm{\sigma},\bm{e},\bm{a},(b_{i}=0,\mathbf{1}{-i}))\right).$
On the other hand, if this inequality is not strict, then the platform $i$ would have an incentive to increase its price. However, if the price is equal to
$\beta\left(\mathcal{I}(\bm{\sigma},\bm{e},\bm{a},\mathbf{1})-\mathcal{I}(\bm{\sigma},\bm{e},\bm{a},(b_{i}=0,\mathbf{1}{-i}))\right),$
platform also has no incentive to deviate. This implies that the prices given in Proposition 1 constitute an equilibrium and that this is the unique price equilibrium. This completes the proof. $\blacksquare$
@@
Suppose the contrary that there exists a platform equilibrium strategy $(\bm{\sigma},\bm{e})$ for which $e_{i}=1$, $a_{i}=0$ and $\bm{a}\in\mathcal{A}^{\mathrm{E}}(\bm{\sigma},\bm{e})$. We argue that platform $i$ has a profitable deviation. In particular, notice that the current utility of platform $i$ is $c_{i}$, as the buyer will not pay any non-zero price for this platform’s data. Now, if platform $i$ chooses $\tilde{e}_{i}=1$, $\tilde{p}_{i}=0$ and $\tilde{\sigma}_{i}=\infty$, then in any $\bm{a}\in\mathcal{A}^{\mathrm{E}}((\tilde{\sigma}_{i},\bm{\sigma}_{-i}),\bm{e})$ we must have that $a_{i}=1$ and therefore the platform $i$’s utility increases to $\mathcal{I}_{i}-c_{i}$. $\blacksquare$
@@
$\mathcal{U}_{1}((\sigma_{1},\sigma_{2})),\mathbf{1},\mathbf{p}^{*},\mathbf{1})=$
$\quad\frac{1}{2}+\beta\left(\mathcal{I}(\sigma_{1},\sigma_{2})-\mathcal{I}(\sigma_{2})\right)\geq\frac{1}{2}+\beta\mathcal{I}(\tilde{\sigma}_{1})=\mathcal{U}_{1}((\tilde{\sigma}_{1},\sigma_{2})),(1,0)),\mathbf{p}^{*},\mathbf{1}).$ (17)
@@
and thus, $\mathcal{I}(\sigma_{1},\sigma_{2})=1/\alpha$. Using this, we can rewrite the condition (A) as
@@
[…]
@@
We can verify that the right-hand side minus the left-hand side is equal to zero at $\bar{\alpha}\approx 1.88$ and positive afterward. Notice that $\eta$ and therefore the condition (26) is symmetric with respect to both platforms, and consequently, considering the other deviation to $(\sigma_{1}^{A},\tilde{\sigma}_{2})$ leads to the exact same condition on $\alpha$. This completes the proof of our claim that both conditions hold if $\alpha\geq\bar{\alpha}$ and neither if $\alpha<\bar{\alpha}$. $\blacksquare$
@@
Proof of Lemma 7: Using Proposition 2, we know that the equilibrium noise variance belongs to $\Sigma_{\mathbf{1},\bm{e}}$. We next argue that it cannot belong to the interior of $\Sigma_{\mathbf{1},\bm{e}}$. To see this, suppose the contrary that $\bm{\sigma}\in\mathrm{int}(\Sigma_{\mathbf{1}},\bm{e})$ is an equilibrium. There exists $i\in[K]$ such that, by locally decreasing $\sigma_{i}$, we remain in the set $\Sigma_{\mathbf{1},\bm{e}}$. We claim that the platform $i$ has an incentive to decrease its noise variance. First, note that by decreasing its noise variance, we remain in the set $\Sigma_{\mathbf{1},\bm{e}}$ and that the user still finds it optimal to share with platform $i$. However, the payment she receives from the buyer, which is given by
@@
$\frac{1}{2}-c_{i}+\beta\left(\mathcal{I}(\bm{\sigma},\bm{e},(a_{i}=1,\bm{a}_{-i}),\mathbf{1})-\mathcal{I}(\bm{\sigma},\bm{e},(a_{i}=0,\bm{a}_{-i}),\mathbf{1})\right)\geq 0.$ (29)
@@
Notice that, for any given $\bm{e}_{-i}$, the denominator is positive because the amount of leaked information strictly decreases when one less user shares information. Finally, $\bar{\beta}$ is finite becasue the outer maximum is taken over finitely many elements.
@@
We next prove that, assuming a set of platforms that have entered the market, for large enough $\alpha$, there exists an equilibrium. Without loss of generality, let us assume that all platforms have entered the market. $\bm{\sigma}\in\Sigma_{\mathbf{1},\mathbf{1}}\cap\Sigma_{\mathbf{0},\mathbf{1}}$ is an equilibrium noise level if no platform has a deviation. Let us consider platform 1. For this platform, the possible deviations are to $\Sigma_{\bm{a},\mathbf{1}}$ where $a_{1}=1$ and $\bm{a}_{-1}$ can be any vector in $\{0,1\}^{K-1}$. We next argue that $\bm{\sigma}\in\Sigma_{\mathbf{1},\mathbf{1}}\cap\Sigma_{\mathbf{0},\mathbf{1}}$ for which
@@
for any $S\subseteq[K]$ denote the leaked information about $\theta$ when the data of the user is shared with the buyer through platforms in the set $S$, and platform $i\in S$ is using noise variance $\sigma_{i}^{2}$.
@@
$\overset{(a)}{=}$ $\bm{\gamma}^{T}\left(\bm{\gamma}\bm{\gamma}^{T}+(\frac{1}{t}-1)\mathrm{diag}(\gamma_{1}^{2},\dots,\gamma_{K}^{2})\right)^{-1}\bm{\gamma}$
$\overset{(a)}{=}$ $\bm{\gamma}^{T}\left(\frac{1}{\frac{1}{t}-1}\mathrm{diag}((1/\gamma_{i}^{2})_{i=1}^{k})-\frac{\frac{1}{\frac{1}{t}-1}\mathrm{diag}((1/\gamma_{i}^{2})_{i=1}^{k})~\bm{\gamma}^{T}\bm{\gamma}~\frac{1}{\frac{1}{t}-1}\mathrm{diag}((1/\gamma_{i}^{2})_{i=1}^{k})}{1+\bm{\gamma}^{T}\frac{1}{\frac{1}{t}-1}\mathrm{diag}((1/\gamma_{i}^{2})_{i=1}^{k})\bm{\gamma}}\right)\bm{\gamma}$
@@
Let us consider the deviation of platform 1 so that the corresponding user equilibrium is $(a_{1}=1,\bm{a}_{-i})$ with $\bm{a}_{-i}$ having i1i-1 ones and KiK-i zeros. Notice that if such a deviation is not possible by increasing $\sigma_{1}$, then this is not a profitable deviation for platform 1, and we now verify that even if such a deviation is possible, it is not profitable for platform 1. Given that $\bm{\sigma}\in\Sigma_{\mathbf{1},\mathbf{1}}\cap\Sigma_{\mathbf{0},\mathbf{1}}$, we see that the user is indifferent between sharing with all platforms and sharing with none of them. Therefore, we have
@@
Using Lemma 9, the above equality gives us the following.
@@
Given the above deviation to $(\tilde{\sigma}_{1},\dots,\sigma_{K})$, the user utility gives us
$\frac{K}{2}-\alpha\mathcal{I}(\tilde{\sigma}_{1},\dots,\sigma_{K})\leq\frac{i}{2}-\alpha\mathcal{I}(\tilde{\sigma}_{1},\dots,\sigma_{i}).$
Substituting the closed-form expressions of Lemma 9 into the above equality gives us the following.
$t^{\prime}\leq\frac{2-4\alpha+K-i+\sqrt{16\alpha^{2}+(K-i)^{2}-8\alpha K}}{2+4\alpha-2K},$ (31)
@@
Now, for this deviation to be not profitable, we must have the following.
@@
Note that the right hand side is increasing in $t^{\prime}$, and therefore, it suffices to show it holds for $t^{\prime}$ given in (31). Again, using the closed-form expressions of Lemma 9 and writing $t^{\prime}$ and $t$ in terms of $\alpha$, the above inequality becomes equivalent to
@@
$\frac{3K-i}{4\alpha}+\frac{1-K}{2\alpha-1}+\frac{i-1}{2\alpha-1+i-K}-1+\frac{\sqrt{16\alpha^{2}+(K-i)^{2}-8\alpha K}}{4\alpha}$ (32)
@@
Notice that for any other set $S\subseteq[K]$ of the platforms that have entered the market, the above bound on $\alpha$ guarantees the existence of a platform equilibrium strategy as $|S|\leq K$. This completes the proof. $\blacksquare$
@@
Clearly, there is no uniform minimum privacy mandate that can make the user better off because banning data sharing maximizes the user’s gain from services and minimizes the privacy loss.
@@
$\frac{1}{2}K-\alpha\mathcal{I}(\underbrace{\bar{\sigma},\dots,\bar{\sigma}}_{K\text{ many}})>\frac{1}{2}(\text{ number of low-cost platforms}).$
For any $\alpha$, as $\bar{\sigma}$ increases, the leaked information goes to zero, and therefore, so long as there exists at least one high-cost platform, a $\bar{\sigma}$ for which the above inequality holds exists. We next argue that all platforms enter the market for a large enough $\beta$. To see this, we need to have
@@
for all $i\in[K]$. Therefore, letting $\beta\geq\tilde{\beta}$ where
@@
Similar to the proof of Proposition 7 and Theorem 1, we choose $\bar{\alpha}$ and $\hat{\beta}$ large enough so that an equilibrium exists and all platforms enter the market. Now, suppose $\bar{\sigma}$ is the optimal uniform minimum privacy mandate so that all platforms enter the market, the user utility is